# Memorization, Emergence, and Explaining Reversal Failures: A Controlled Study of Relational Semantics in LLMs

### v1

@@ A.3.3 SFT Corpus Construction
(7) (Michael Brett Ferguson, entertainment manager)
@@
(5) Gabriel Bode Arnold is employed as an it consultant.
@@ A.4.1 Effect of Increasing Templates Under Fixed Train Samples
A.4.2 More Detail and Experiment of Layer-wise Analysis
@@ A.4.3 Effect of SFT Data Fraction
To better understand the role of SFT, we conduct an additional experiment that varies the amount of SFT data. Our SFT corpus is derived from the Train Content: for each sentence verbalized from a triple, we construct a corresponding QA pair to equip the pre-trained model with question-answering capability. In all previous experiments, we fine-tuned on the full SFT corpus. Since the model pre-trained with $N=2000$ and $K=10$ already achieves strong QA performance after fine-tuning, we fix this pre-training setting and run SFT using only $10\%$ to $100\%$ of the SFT corpus. As shown in Figure 11, we evaluate $\text{Q}_{\text{non}}^{\text{People}}$ (forward and reverse), $\text{Q}_{\text{non}}^{\text{Job}}$ (reverse), and $\text{Q}_{\text{Logic}}$ (Inversion and Symmetry), and additionally test QA on training content excluded from SFT (e.g., when using 10% of the SFT corpus, we sample QA instances from the remaining 90% portion). We find that performance is modest at 10%, but starting from 20% the accuracies exhibit a clear turning point and already reach high values, indicating that even a small fraction of the SFT corpus is sufficient for effective fine-tuning.
@@ A.5.1 Implementation Details
  • Bidirectional Attention: We remove the causal masking in the self-attention layers. This allows every token to attend to the entire sequence context (both prefix and suffix) simultaneously, which is fundamental for learning joint probability distributions.
  • Training Objective: The model is trained using a Bernoulli masking strategy. At each training step, tokens are independently masked with a probability $t\sim U(0,1)$, and the model optimizes the re-weighted cross-entropy loss on the masked positions.
  • Inference Strategy: We employ a block-wise generation strategy with low-confidence re-masking. Specifically, the answer is generated in blocks of 4 tokens. Within each block, we iteratively refine the tokens by re-masking positions with the lowest confidence scores, allowing the model to self-correct based on the global bidirectional context.
@@ A.5.2 Data Adjustment
  • Raw (25%): Original complete paragraphs to maintain document-level coherence.
  • Single (25%): Randomly sampled individual sentences. This forces the model to learn atomic syntax and entity structures without reliance on external context.
  • Subset (50%): Random subsets containing 2 to N1N-1 sentences. This component is crucial as it simulates “missing link” scenarios, compelling the model to internalize the underlying logical rules to recover the missing information during training.
@@ A.5.4 Discussion: Why does Reverse outperform Forward?
(3) {person_a} acts in the role of {relationship} to {person_b}, {person_a} holds the relation of {reverse_relationship} to
@@
(2) Noah Dylan Martinez is the wife of Gregory Joel Henderson
@@
(4) Noah Dylan Martinez is the friend of Raymond Clayton Morris

### v2

@@ A.3.3 SFT corpus construction
(7) (Michael Brett Ferguson, job, entertainment manager)
@@
(5) Gabriel Bode Arnold is employed as a consultant.
@@ A.4.1 Effect of increasing templates under fixed train samples
A.4.2 More details and experiments on layer-wise analysis
@@ A.4.3 Effect of SFT data fraction
To better understand the role of SFT, we conduct an additional experiment that varies the amount of SFT data. Our SFT corpus is derived from the Train Content: for each sentence verbalized from a triple, we construct a corresponding QA pair to equip the pre-trained model with question-answering capability. In all previous experiments, we fine-tuned on the full SFT corpus. Since the model pre-trained with $N=2000$ and $K=10$ already achieves strong QA performance after fine-tuning, we fix this pre-training setting and run SFT using only $10\%$ to $100\%$ of the SFT corpus. As shown in Figure 11, we evaluate $\text{Q}_{\text{Mem}}^{\text{People}}$ (forward and reverse), $\text{Q}_{\text{Mem}}^{\text{Job}}$ (reverse), and $\text{Q}_{\text{Logic}}$ (Inversion and Symmetry), and additionally test QA on training content excluded from SFT (e.g., when using 10% of the SFT corpus, we sample QA instances from the remaining 90% portion). We find that performance is modest at 10%, but starting from 20% the accuracies exhibit a clear turning point and already reach high values, indicating that even a small fraction of the SFT corpus is sufficient for effective fine-tuning.
@@ A.5.1 Implementation details
  • Bidirectional Attention. We remove the causal masking in the self-attention layers. This allows every token to attend to the entire sequence context (both prefix and suffix) simultaneously, which is fundamental for learning joint probability distributions.
  • Training Objective. The model is trained using a Bernoulli masking strategy. At each training step, tokens are independently masked with a probability $t\sim U(0,1)$, and the model optimizes the re-weighted cross-entropy loss on the masked positions.
  • Inference Strategy. We employ a block-wise generation strategy with low-confidence re-masking. Specifically, the answer is generated in blocks of 4 tokens. Within each block, we iteratively refine the tokens by re-masking positions with the lowest confidence scores, allowing the model to self-correct based on the global bidirectional context.
@@ A.5.2 Data adjustment
  • Raw (25%). Original complete paragraphs to maintain document-level coherence.
  • Single (25%). Randomly sampled individual sentences. This forces the model to learn atomic syntax and entity structures without reliance on external context.
  • Subset (50%). Random subsets containing 2 to N1N-1 sentences. This component is crucial as it simulates “missing link” scenarios, compelling the model to internalize the underlying logical rules to recover the missing information during training.
@@ A.5.4 Discussion: why does reverse outperform forward?
(3) {person_a} acts in the role of {relationship} to {person_b}, {person_b} holds the relation of {reverse_relationship} to
@@
(2) Noah Dylan Martinez is the wife of Gregory Joel Henderson.
@@
(4) Noah Dylan Martinez is the friend of Raymond Clayton Morris.